MIDC News
Year of 2009
December 30, 2009
- 15:59 MST: The SDM-INT8 used at NWTC M2 for measuring wind speed was removed at 12:35 MST today. The wind speed sensors are now measured by the 4 pulse and 2 control ports on the CR23X data logger. -Afshin Andreas
- 12:24 MST: The SRRL YES Total Sky Imager has been down since sunrise. Instrument will be troubleshot later. -Afshin Andreas
- 09:58 MST: The SRRL BMS Instantaneous data stream for the radiometer tower was changed at 9:36 MST. Downwelling CM3, Upwelling CM3, Atmos. Net IR was replaced with Global 40-South PSP, CNR1 case temp, and Instrument Net DW-CG3. Note: This data stream is for internal use only. -Afshin Andreas
- 15:05 MST: The SRRL BMS Downwelling and Upwelling PIRs were swapped out with calibrated instruments today from 14:30 to 14:45 MST. Logger was updated with new coefficients at 14:41 MST. -Afshin Andreas
- 13:22 MST: There was a power outage (and bad power) at NREL on December 9, 2009 @ 10:52 MST; SRRL BMS downtimes varied and are as follows:
- 12/9/09 10:52 to 14:22 MST: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210, SPN1, TSI-880, POM-01, Snow Depth.
- 12/9/09 12:11 to 14:22 MST: Direct CH1, NIP 1+2, LI-201, NIP RG780, Diffuse 8-48, CM22, PSP, CUV4.
- 12/9/09 11:50 to 14:25 MST: SRRL PGS-100 and SkyCam.
- 12/9/09 11:00 to 12/10/09 9:00 MST: SRRL LI-1800/Global-40.
- 12/9/09 12:05 to 12/10/09 9:41 MST: OTF MS-700/Global-40.
- 12/9/09 10:00 to 12/17/09 12:15 MST: SRRL TSR-1s.
- 08:42 MST: The SRRL BMS Instrument Pictures page was updated yesterday with new instrumentation. However, the two Sci-Tec trackers still need to be updated to reflect the latest changes to BMS. -Afshin Andreas
- 12:25 MST: May 2009 through November 2009 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 10:42 MST: August 2009 through November 2009 Monthly Data Sets & Reports for BSC have been generated. -Afshin Andreas
- 10:42 MST: May 2009 through November 2009 Monthly Data Sets & Reports for ECSU have been generated. -Afshin Andreas
- 14:11 MST: The RG780 filter for the SRRL BMS Global RG780 PSP was replaced today at approx 14:00 MST. -Pete Gotseff
- 16:16 MST: The SRRL Direct Normal PGS-100 Spectroradiometer was calibrated today using outdoor spectral data measured by a NIST traceable LICOR LI-1800. New PGS-100 calibration will go into affect on 11/18/2009 at sunrise. -Afshin Andreas
- 14:45 MST: The SRRL BMS HMP-45C Temp/RH probe (tower) was swapped out with a calibrated sensor @ 14:14 through 14:15 MST today. -Afshin Andreas
- 13:54 MST: The SRRL PGS-100 Direct Normal Spectroradiometer was returned to service @ 13:50 MST today. The shutter solenoid was replaced at the factory. -Afshin Andreas
- 11:19 MST: All SRRL BMS UV instruments were calibrated on 10-26-2009, this includes: UV-S-B-T, CUVB1, MS210W, UVB-1, 501A, TUVR, CUV4 (both), UV-S-A-T, CUVA1, CUVA2, CUVB2. Logger was updated with new cal factors today @ 10:42 MST. -Afshin Andreas
- 10:00 MST: The imbedded linux box inside the SRRL TSI-880 is malfunctioning. Cloud cover is unavailable. -Afshin Andreas
- 07:21 MST: The shutter on the The SRRL PGS-100 Direct Normal Spectroradiometer is sticking again (zero readings for all wavelengths). The unit will be removed and shipping back to PREDE/Japan. -Afshin Andreas
- 12:09 MST: Three Yankee Environmental Systems (YES) Thermopile Shadowband Radiometer (TSR-1) instruments have been installed at SRRL. Data is available since August 28, 2009. -Afshin Andreas
- 09:54 MST: The SRRL BMS Global CM3 (on CNR1 at Radiometer tower) was calibrated using BORCAL 2009-04 reference data. New cal factor was entered in the logger @ 9:51 MST today. -Afshin Andreas
- 13:25 MST: The SRRL PGS-100 Direct Normal Spectroradiometer was returned to service @ 10:05 MST today. Unit was inspected, but no problem was found with shutter. Instrument appears to be working now. -Afshin Andreas
- 13:23 MST: There was a planned power outage at SRRL (and OTF) on Saturday September 12, 2009; SRRL BMS downtimes varied and are as follows:
- 9/12/09 06:13 to 10:37 MST: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210, SPN1, TSI-880.
- 9/12/09 7:15 to 10:55 MST: LI-1800 Global 40 Spectroradiometer, SkyCam.
- 9/11/09 15:52 to 9/14/09 9:03 MST: OTF MS-700 Global 40 Spectroradiometer.
-Afshin Andreas
- 09:20 MST: The NWTC M2 wind speed, wind direction, and temperature sensors were swapped out with calibrated sensors, yesterday from 8:45 to 11:50 MST. The multipliers and offsets for the wind speed sensors and the coefficients for the temperature sensors were updated in the logger at 9:14 MST today. -Afshin Andreas
- 13:18 MST: The SRRL PGS-100 Direct Normal Spectroradiometer was removed from service at 07:50 MST today. Instrument has an intermittent problem with the shutter sticking in the closed position, which is causing zero irradiance readings. -Afshin Andreas
- 15:15 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service today @ 12:17 MST, calibrated against NIST F620, and returned to service @ 14:28 MST. -Afshin Andreas
- 15:03 MST: The NREL OTF MS-700 spectroradiometer was returned to service @ 15:00 MST today. -Afshin Andreas
- 12:42 MST: The SRRL BMS UV-S-B-T was removed from service at 9:45 MST today, due to a shattered dome. -Afshin Andreas
- 14:03 MST: Data Lab (indoor) Temperature and Relative Humidity was added to SRRL BMS today @ 13:19 MST. -Afshin Andreas
- 07:34 MST: The NREL OTF MS-700 spectroradiometer was removed from service for calibration on July 29, 2009 @ 10:10 MST. Unit was calibrated against NIST F620 on August 11 and will be returned to service soon. -Afshin Andreas
- 11:49 MST: There is a problem with the SRRL BMS Shadowband PSP, which started today @ 9:30 MST. Signal is high and somewhat erratic. -Afshin Andreas
- 15:10 MST: The following SRRL BMS instruments were swapped out today between appox 14:24 to 14:43 MST with instruments that were calibrated in BORCAL 2009-01/02: Global 90-North/East/South/West PSPs. Cal factors were updated in logger at 14:43 MST. -Afshin Andreas
- 09:50 MST: There is a problem with the SRRL BMS Global SP-110. Instrument has been reading low since July 8. SP-100 channel has been shorted for troublshooting. -Afshin Andreas
- 08:55 MST: There has been intermittent problems with the SRRL PGS-100 Direct Normal Spectroradiometer since July 8 @ 13:40 MST. Many of the scans are reading zero for all wavelengths. -Afshin Andreas
- 15:12 MST: The following SRRL BMS instruments were swapped out today between appox 14:30 to 14:40 MST with instruments that were calibrated in BORCAL 2009-01/02: 40-South & Upwelling PSPs, Global Normal CM-21, and 90-North/East/West LI-200s. Cal factors were updated in logger at 14:36 MST. -Afshin Andreas
- 11:01 MST: Data from the Solar Resource & Meteorological Assessment Project (SOLRMAP): La Ola Lanai, Hawaii - Rotating Shadowband Radiometer (RSR) is now available on MIDC. -Afshin Andreas
- 07:30 MST: The following SRRL BMS instruments were swapped out yesterday between 14:43 to 14:49 MST with instruments that were calibrated in BORCAL 2009-01/02/03: Global PSP RG780, Direct NIP RG780, Zebra PSP, Diffuse PSP Shadowband, Diffuse PSP, & Diffuse 8-48. Cal factors for these instruments as well as the Global CM22 and Global SP-110 were updated in the logger at 14:50 MST. -Afshin Andreas
- 13:07 MST: The Public MIDC Server was upgraded to a new faster server on July 10. However the ZIP downloads from daily data, site comparison, and solar position are not working at this time. Problem is being investigated. -Afshin Andreas
- 07:31 MST: There is a problem with the SRRL BMS Global 90-East LI-200, it has been reading zero since 7/19/2009 at 07:16 MST. -Afshin Andreas
- 10:03 MST: The following SRRL BMS instruments were calibrated in BORCAL 2009-01 & 2009-02: Global Horizontals: CM6b, PSP (ventilated), PSP, TSP-700, TSP-1, SP-Lite, LI-200, SPN1; Direct Normals: CH1, NIP #2, NIP #1, LI-201; Diffuse CM22, SPN1. Logger program was updated at 9:26 MST today with new cal factors. Responsivity Coefficients for Global PSPs were also updated at 9:26 MST today. -Afshin Andreas
- 14:45 MST: Fixed a bug with daily total irradiance calculations from raw data for sites that do not collect data at night (applies to SRRL AOCS and pre-February-1993 SRRL ATI data only), which caused incorrect totals. Note: This required the ATI data stream to be re-configured into 4 time frames (instead of 1), due to data logger output intervals changing throughout it's life time. -Afshin Andreas
- 13:55 MST: The following SRRL BMS LI-200 pyranometers were swapped out today with instruments that were calibrated in BORCAL 2009-01: 40-South, 90-South, and Upwelling. Logger was updated with new cal factors on today @ 13:40 MST. -Afshin Andreas
- 10:33 MST: July 2008 Monthly Data Sets & Reports for BSC have been generated. -Afshin Andreas
- 10:32 MST: July 2008 through April 2009 Monthly Data Sets & Reports for ECSU have been generated. -Afshin Andreas
- 08:55 MST: July 2008 through April 2009 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 09:08 MST: The heater for the SRRL BMS Net Radiometer CNR1 was turned off at 9:07 MST today. The heater will no longer be used. -Afshin Andreas
- 12:36 MST: The SRRL BMS Downwelling CG3 (s/n 010284) was calibrated on 3/23/2009, new coefficients were entered into the logger today @ 12:30 MST. -Afshin Andreas
- 10:43 MST: The NWTC M2 Global PSP was swapped from s/n 31555F3 to s/n 20517F3 at 08:57 through 09:06 MST today. New calibration factor was entered in logger today @ 10:39 MST. -Afshin Andreas
- 10:26 MST: SRRL BMS data collection resumed late afternoon on April 20. The SC32A RS-232 Optical Isolator which failed was replaced. There was no loss of data. -Afshin Andreas
- 09:50 MST: There is a problem with the SRRL BMS CR23X datalogger, last 1-minute data collection occured on April 17, 2009 at 14:55 MST. Logger is non-responsive, problem will be troubleshoot during normal business hours. -Afshin Andreas
- 13:25 MST: The SRRL BMS Downwelling PIR (s/n 31192F3) and CG4 (s/n 010548) were calibrated on 3/23/2009, new coefficients were entered into the logger today @ 13:21 MST. -Afshin Andreas
- 14:14 MST: The SRRL Direct Normal PGS-100 Spectroradiometer was calibrated on 4/10/2009 using data from 4/8/2009 measured by a NIST traceable LICOR LI-1800. -Afshin Andreas
- 07:14 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service yesterday @ 9:07 MST, calibrated against NIST F620, and returned to service @ 9:33 MST yesterday. -Afshin Andreas
- 08:40 MST: Data from the Sacramento Municipal Utility District (Anatolia) RSR is now available on the MIDC. Data collection started Februrary 3, 2009. -Afshin Andreas
- 09:23 MST: There has been an unknown problem with the MIDC public server for the past month. The server is very slow, downloads being the worst. NREL users should be using the internal MIDC server for optimal perfomance. -Afshin Andreas
- 10:57 MST: The NREL/OTF EKO MS-700 Spectroradiometer was returned to service today @ 9:31 MST. -Afshin Andreas
- 15:56 MST: The NREL/OTF EKO MS-700 Spectroradiometer was removed from service on Feb 3 @ 9:09 MST for calibration. Unit was calibrated today and will be returned to service soon. -Afshin Andreas
- 15:14 MST: The datalogger at Humboldt State University (HSU) is now on a network interface and data is automatically collected by MIDC. Data is currently updated every hour. -Afshin Andreas
- 13:30 MST: The SRRL BSRN proto-type station has been permanently shutdown as of 11:00 MST today. -Afshin Andreas
- 13:41 MST: Updated the MIDC SPA Calculator with the latest SPA C source code, which allows the timezone to go up to +/-18 hours. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]